the length of a rectangle is 9 centimeters more than half the width. find the length if the perimeter is 60 centimeters
Found 2 solutions by mananth, benazir.sj@gmail.com:
Answer by mananth(16949)

let width be x cm
length = x/2 +9
..
Perimeter = 2(L+W)
=2(x+(x/2)+9)
=2(3x/2 +9)
=2(3x+18)/2
=3x+18
..
3x+18=60
add -18
3x+18-18=60-18
3x=42
divide by 3
x= 14 the width
..
length = x/2 +9
14/2 +9
=7+9
=16cm length
..
CHECK
2*(14+16)=60

Answer by benazir.sj@gmail.com(109)

Let the width of the rectangle be x
and the length is 9 centimeters more than half the width
that is
length = 9+x/2
given perimeter =60cm
perimeter =2(l+w)
60 =2(9+x/2+x)
60=2(9+3x/2)
60 =18+2*3x/2
60 =18+3x
60-18=3x
42=3x
x=14
width =14cm
length =9+x/2
=9+14/2
=9+7
=16cm
length=16cm and width =14cm